OpenStudy (anonymous):

b^3-125. Use the sum of two cubes or the difference of two cubes pattern to factor the following. I have (b-5)(b^2+5b+25) But I don't know how to get there. Someone just gave me the answer. Anyone willing to tackle showing me how to arrive at the answer? Please and thank you.

OpenStudy (anonymous):

\[(b^3-125) = (b-5)(b^2+5b+25)\]

OpenStudy (anonymous):

The formula I used is \[(a^3-b^3) = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

first you're writing 125 as 5^3 then=>b^3-5^3=(b-5)(b^2+5b+25)

OpenStudy (anonymous):

Thank you both, I appreciate the input!